# Information scrambling in quantum-walks: discrete-time formulation of Krylov complexity

### v2

@@ Abstract
We study information scrambling — a spread of initially localized quantum information into the system’s many degree of freedom — in discrete-time quantum walks. We consider out-of-time-ordered correlators (OTOC) and K-complexity as probe of information scrambling. The OTOC for local spin operators in all directions has a light-cone structure which is “shell-like”. As the wavefront passes, the OTOC approaches to zero in the long-time limit, showing no signature of scrambling. The introduction of spatial or temporal disorder changes the shape of the light-cone akin to localization of wavefuction. We formulate the K-complexity in system with discrete-time evolution, and show that it grows linearly in discrete-time quantum walk. The presence of disorder modifies this growth to sub-linear. Our study present interesting case to explore many-body phenomenon in discrete-time quantum walk using scrambling.
@@ I Introduction
There remain ambiguity is describing the process of quantum scrambling. One method to probe it involves using an out-of-time-ordered correlator [17, 1]. For systems that exhibit a semi-classical limit or have a large number of local degrees of freedom, this correlator shows exponential growth, which can be used to identify a quantum counterpart to the Lyapunov exponent (LE), thus linking it to classical chaos [18]. Another approach involves studying the evolution dynamics of operators in Krylov space. Here, operator growth is measured by the “K-complexity”, indicating the extent of delocalization of initial local operators evolving under Heisenberg evolution under the system Hamiltonian [19, 20, 21]. It is speculated that this K-complexity grows exponentially in most generic nonintegrable systems [19]. This exponential growth in K-complexity can be used to extract LE, establishing a connection with out-of-time-ordered correlators [6]. Recent studies have explored K-complexity in various systems such as Ising models [22, 23, 24], Sachdev-Ye-Kitaev (SYK) models [25, 26, 27], quantum field theories [28, 29, 30, 31, 32, 33], many-body localization system [34, 35], and open quantum systems [36, 37, 38, 39, 40].
@@
Experimentally tunable toy models serve as valuable tools for investigating different phenomena from theoretical physics. One such tool, which we will pursue in this work, is the quantum walk — a quantum version of the classical random walk [41, 42]. Specifically, we examine discrete-time quantum walks, which have been previously employed to simulate controlled dynamics in quantum systems [43, 44, 45] and to construct quantum algorithms [42]. These quantum walks are implemented experimentally using both lattice-based quantum systems and circuit-based quantum processors [46, 47, 48]. The adaptability of quantum walks, allowing for the experimental modeling of various phenomena like topological effects [45, 49], therefore, also positions them as a promising platform for studying scrambling.
Figure 1: Schematic showing the “shell-like” structure of OTOC in which as the wavefront passes the operator site $l$, the OTOC approaches to zero in the long-time limit. The horizonal line shows different time slices increasing in upward direction. We note that the site where OTOC is non-zero at initial time, goes to zero at late time, implying that operator has no support on site.
@@
In this article, we study the out-of-time-ordered correlator (OTOC) and K-complexity for different operators in the exactly-solvable one-dimensional discrete-time quantum walk. Our study of OTOC for different spin operators shows universal “shell-like” structure so as the wavefront passes, the OTOC goes to zero in the long-time limit, in other words, operator has no support on the site, implying absence of scrambling (See Fig. 1). On the other hand, we show that K-complexity grows linearly in time, akin to approximate orthonorgonality of operator at each time-step. We further study the effect of disorder which generally results in slowdown of information scrambling. In both spatial and temporal disorder, the shape of light-cone deforms showing no scrambling beyond localization length. The K-complexity growth transit from linear to sub-linear showing saturation at late-times reflecting the localization of operator.
@@ II.1 Out-of-time-ordered correlators
Out-of-time-ordered correlators (OTOCs) provide a means to quantify the evolution of operators. Let’s consider two local operators, $W$ and $V$, within a one-dimensional spin chain. The idea is to probe the spread of $W(t)=e^{iHt}We^{-iHt}$ using another operator $V$, typically a simple spin operator positioned at a distance $l$ from $W$ evolving under system Hamiltonian $H$. To do this, one considers the expectation value of the squared commutator,
@@ II.2.1 Continuous-time evolution
where $\mathcal{L}$ is Hermitian Liouvillian superoperator given by $\mathcal{L}=[H,\,\bullet\,]$. Therefore, the operator $\mathcal{O}(t)$ can be written as a span of the nested commutators with the initial operator i.e. $\{\mathcal{L}^{n}\mathcal{O}_{0}\}_{n=0}^{\infty}$. A orthonormal basis $\{|O_{n})\}_{n=0}^{K-1}$ can be constructed from this nested span of commutators, by choosing a certain scalar product (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ) on operator space using a form of Gram-Schmidt orthogonalization known as Lanczos algorithm. The dimension of Krylov space $K$ obeys a bound $K\leq D^{2}-D+1$, where $D$ is the dimension of the state Hilbert space [21]. In the krylov basis $\{|O_{n})\}$, the Liouvillian takes the tridiagonal form $\mathcal{L}|O_{n})=b_{n+1}|O_{n+1})+b_{n}|O_{n})$, where $b_{n}$ are known as Lanczos coefficients that are tied to chaotic nature of the system at hand [23]. We can write the expansion of the operator $\mathcal{O}(t)$ in terms of constructed krylov basis as
@@
The amplitudes $\phi_{n}(t)$ evolve according to the recursion relation $\dot{\phi}_{n}(t)=b_{n-1}\phi_{n-1}(t)-b_{n}\phi_{n+1}(t)$ with the initial conditions $\phi_{n}(0)=\delta_{n,0}$ The recursion relation suggests that the Lanczos coefficients $b_{n}$ are hopping amplitudes for the initial operator $\mathcal{O}_{0}$ localized at the initial site to explore the Krylov chain. With time, the operator gains support away from the origin in Krylov chain reflects the growth of complexity as higher Krylov basis vectors are required in operator expansion. To quantify this, one defines the average position of the operator in Krylov chain — called the Krylov complexity as
@@
where $\mathcal{K}=\sum_{n=0}^{K-1}n|O_{n})(O_{n}|$ is position operator in the Krylov chain. For our purpose, we will use the infinite-temperature inner product, also known as Frobenius inner product :
@@ II.2.2 Discrete-time evolution
To formulate the K-complexity for system with discrete-time evolution such that $\mathcal{O}_{n}=U^{\dagger}_{n}\mathcal{O}_{n-1}U_{n}=\mathcal{U}_{n}[\mathcal{O}_{n}]$, where $n=1,2,\ldots$, $U_{n}$ is the unitary operator which describes the evolution of the system at time-step $t=nT$ with step-size $T$, and $\mathcal{U}_{n}$ is unitary superoperator given by $\mathcal{U}_{n}=U_{n}^{\dagger}\bullet U_{n}$. We define the Krylov basis by choosing $|O_{0})=\mathcal{O}_{0}$ and then recursively orthogonalizing each $\mathcal{O}_{n}$ with all the $|O_{n})$ for $i<n$. This choice of basis (represented by $\mathbb{O}$) maximizes the cost function defined as [59]
@@
with respect to arbitrary choice of orthonormal set $\mathbb{B}=\{|B_{i}):i=0,1,2,\ldots\}$. The coefficients $\zeta_{i}$ (refer to as “weight function”) are positive, increasing sequence of real numbers. The optimal choice of the basis follows from induction method. We fix the initial state as the first state of Krylov basis $\mathcal{O}_{0}=|O_{0})$. Assume the first $N$ vectors of certain basis $\mathbb{B}$ are same as the Krylov basis, i.e., $|B_{i}\rangle=|O_{i}\rangle$ for $i=0,1,\ldots,N-1$. By assumption
@@
Therefore, the cost associated with two basis are same for $n\leq N-1$. The next state $|O_{N}\rangle$ into a part belonging to the Krylov subspace $|O_{i}\rangle$, for $i\leq N-1$, and a part perpendicular to it i.e.
@@
Since the evolution operator itself is used as generator of Krylov basis, we can bound the maximum possible growth of K-complexity. For any system evolving unitarily in discrete-time, the K-complexity can grow at most linearly with time $t$. To prove this, we consider maximizing the complexity at anytime $t$ with respect to expansion coefficients, i.e.,
@@
with constraint $\sum_{i}|\phi_{i,n}|^{2}=1$. From the orthonormalization construction, it follows that the expansion coefficients $\phi_{i,n}=0\ \forall\ i>n$. Therefore, it follows that the complexity can at most grows linearly with time $t$ corresponding to case where $\phi_{i,n}=\delta_{i,n}$. In terms of operator, it means that $(O_{n^{\prime}}|O_{n})\ \forall\ n^{\prime}<n$ which corresponds to maximally ergodic regime universal in chaotic systems [60].
In the case where the evolution operator is time-independent, we can define a quasi-Hamiltonian $H_{q}$ such that $U=e^{-iH_{q}T}$. In this case, the unitary operator $U$ takes upper Hessenberg form in Krylov basis [60], given by
@@
with $a_{0}=c_{0}$. The quasi-Hamiltonian $H_{q}$ can itself be used in continuous-time setting to define complexity using the notion discussed in Sec. II.2.1. To investigate the relation between the complexities, we consider the following expansion
@@
where $\mathcal{L}_{q}$ is Liouvillian superoperator associated with the Hamiltonian $H_{q}$. The Taylor series is truncated at order $K_{c}$ which depends on the operator $\mathcal{L}_{q}$, step-size $T$ and error tolerance [61]. From Equation (12), it follows that the operator at first time-step $\mathcal{O}_{1}$ will be linear combination of operators $\{\mathcal{L}^{k}_{q}\mathcal{O}_{0}\ |\ k=0,1,\ldots,K_{c}\}$. In other words, the first Krylov basis vector
@@
where the set $\mathbb{O}^{(q)}=\{|O^{(q)}_{k})\ |\ k=0,1,\ldots\}$ corresponds to Krylov basis associated with Liouvillian $\mathcal{L}_{q}$. In general, we will assume that there exist cutoff value $K_{c}^{(i)}$ such that
@@ III Discrete-time quantum walks
corresponding to parameters $(0,\theta,0,3\pi/2)$ which previously has been studied in context of Dirac dynamic [62]. In momentum basis, the unitary operator can be diagonalized to obtain the dispersion relation in space-time continuum limit [63]
@@
It’s important to note that group velocity is maximum (equals to $1$) at $\theta=0$ for all $k$ corresponds to identity as coin operator. The coin parameter $\theta$ controls the variance $\sigma^{2}$ of the probability distribution in the position space and this distribution spreads quadratically faster ($\sigma^{2}\approx[1-t^{2}\sin\theta])$ in position space when compared to the classical random walk [64].
@@ III.1 Disordered discrete-time quantum walk
The mean group velocity drops to zero faster for a walk with spatial disorder resulting in strong localization compared to temporal disorder which leads to weak localization. the localization length is usually a function of the coin parameter $\theta$ given as $\zeta=[\ln\cos\theta]^{-1}$. Both of these disorders have been studied extensively in enhancing the entanglement and non-Markovianity generated between the internal and external degrees of freedom [71, 70].
@@ IV.1 Out-of-time-ordered correlator
In integrable systems such as quantum spin systems [72], one can analytically find the squared commutator to exactly describe the nature of the wavefront and decay. And, although, quantum walk is also an integrable model, it turns out to be particularly hard to find out exact form of squared commutator. This is because the initial operators are local in position space, and the inverse fourier transform becomes hard due to the complexity of functions. However, in case of the Hadamard quantum walk ($\theta=\pi/4$), we can use the known asymptotic results of wavefunction to vaguely argue the nature of decay and wavefront [73]. To this end, the asymptotics for the wave function can be described in three regions using $\alpha=l/t$. The wave function is essentially uniformly spread over the interval between $\mp 1/\sqrt{2}=\mp v_{B}$ where its gross behaviour is like $1/\sqrt{t}$. Outside this interval, the wave function dies out very faster than any inverse polynomial in $t$. At the “wavefront” (±$\pm v_{B}$), there are two peaks of width $\mathcal{O}(t^{1/3})$, where the wave function goes as $t^{-1/3}$. If we approximately translate these wave function behaviour to operator $W^{\mu}_{l}(t)$, we conclude that the squared commutator decays as $\sim 1/t^{2}$ in the long-time limit and goes as $t^{-4/3}$ at the wavefront.
@@ IV.2 K-complexity
where the blue dot represent the presence of term $P_{ij}=|i\rangle\langle j|$ and multiple presence of blue dot represents the linear combination of these terms in operator expansion. The inner product between operators $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$ corresponds depends on the expansion term corresponding to where dot color matches. The inner product between the operators in Eq. (32) is zero i.e. the operators are orthonormal to each other since there are no common blue dots. At $t=2,3$, the operators given by

### v3

# Information scrambling in quantum walks: Discrete-time formulation of Krylov complexity

@@ Abstract
We study information scrambling — a spread of initially localized quantum information into the system’s many degree of freedom — in discrete-time quantum walks. We consider out-of-time-ordered correlators (OTOC) and K-complexity as a probe of information scrambling. The OTOC for local spin operators in all directions has a light-cone structure which is “shell-like”. As the wavefront passes, the OTOC approaches to zero in the long-time limit, showing no signature of scrambling. The introduction of spatial or temporal disorder changes the shape of the light-cone localization of wavefunction. We formulate the K-complexity in system with discrete-time evolution, and show that it grows linearly in discrete-time quantum walk. The presence of disorder modifies this growth to sub-linear. Our study present interesting case to explore many-body phenomenon in a discrete-time quantum walk using scrambling.
@@ I Introduction
There remain ambiguity is describing the process of quantum scrambling. One method to probe it involves using an out-of-time-ordered correlator [17, 1]. For systems that exhibit a semi-classical limit or have a large number of local degrees of freedom, this correlator shows exponential growth, which can be used to identify a quantum counterpart to the Lyapunov exponent (LE), thus linking it to classical chaos [18]. Another approach involves studying the evolution dynamics of operators in Krylov space. Here, operator growth is measured by the “K-complexity”, indicating the extent of delocalization of initial local operators evolving under Heisenberg evolution under the system Hamiltonian [19, 20, 21]. It is speculated that this K-complexity grows exponentially in most generic nonintegrable systems [19]. This exponential growth in K-complexity can be used to extract LE, establishing a connection with out-of-time-ordered correlators [6]. Recent studies have explored K-complexity in various systems such as Ising models [22, 23, 24], Sachdev-Ye-Kitaev (SYK) models [25, 26, 27], quantum field theories [28, 29, 30, 31, 32, 33], the many-body localization system [34, 35], and open quantum systems [36, 37, 38, 39, 40].
@@
Experimentally tunable toy models serve as valuable tools for investigating different phenomena from theoretical physics. One such tool, which we will pursue in this work, is the quantum walk — a quantum version of the classical random walk [41, 42]. Specifically, we examine discrete-time quantum walks, which were previously employed to simulate controlled dynamics in quantum systems [43, 44, 45] and to construct quantum algorithms [42]. These quantum walks are implemented experimentally using both lattice-based quantum systems and circuit-based quantum processors [46, 47, 48]. The adaptability of quantum walks, allowing for the experimental modeling of various phenomena like topological effects [45, 49], therefore, also positions them as a promising platform for studying scrambling.
@@
In this article, we study the out-of-time-ordered correlator (OTOC) and K-complexity for different operators in the exactly solvable one-dimensional discrete-time quantum walk. Our study of OTOC for different spin operators shows a universal “shell-like” structure so as the wavefront passes, the OTOC goes to zero in the long-time limit, in other words, the operator has no support on the site, implying the absence of scrambling (see Fig. 1). On the other hand, we show that K-complexity grows linearly in time, akin to the approximate orthonorgonality of operator at each time-step. We further study the effect of disorder which generally results in the slowdown of information scrambling. In both spatial and temporal disorder, the shape of the light-cone deforms showing no scrambling beyond localization length. The K-complexity growth transititions from linear to sub-linear showing saturation at late times, thus reflecting the localization of the operator.
Figure 1: Schematic showing the “shell-like” structure of OTOC in which as the wavefront passes the operator site $l$, the OTOC approaches to zero in the long-time limit. The horizonal line shows different time slices increasing in an upward direction. We note that the site where OTOC is non-zero at initial time, goes to zero at late time, implying that the operator has no support on site.
@@ II.1 Out-of-time-ordered correlators
OTOCs provide a means to quantify the evolution of operators. Let’s consider two local operators, $W$ and $V$, within a one-dimensional spin chain. The idea is to probe the spread of $W(t)=e^{iHt}We^{-iHt}$ using another operator $V$, typically a simple spin operator positioned at a distance $l$ from $W$ evolving under system Hamiltonian $H$. To do this, we considers the expectation value of the squared commutator,
@@ II.2.1 Continuous-time evolution
where $\mathcal{L}$ is the Hermitian Liouvillian superoperator given by $\mathcal{L}=[H,\,\bullet\,]$. Therefore, the operator $\mathcal{O}(t)$ can be written as a span of the nested commutators with the initial operator i.e. $\{\mathcal{L}^{n}\mathcal{O}_{0}\}_{n=0}^{\infty}$. An orthonormal basis $\{|O_{n})\}_{n=0}^{K-1}$ can be constructed from this nested span of commutators, by choosing a certain scalar product (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ) on operator space using a form of Gram-Schmidt orthogonalization known as the Lanczos algorithm. The dimension of Krylov space $K$ obeys a bound $K\leq D^{2}-D+1$, where $D$ is the dimension of the state Hilbert space [21]. In the Krylov basis $\{|O_{n})\}$, the Liouvillian takes the tridiagonal form $\mathcal{L}|O_{n})=b_{n+1}|O_{n+1})+b_{n}|O_{n})$, where $b_{n}$ are known as the Lanczos coefficients that are tied to chaotic nature of the system at hand [23]. We can write the expansion of the operator $\mathcal{O}(t)$ in terms of the constructed Krylov basis as
@@
The amplitudes $\phi_{n}(t)$ evolve according to the recursion relation $\dot{\phi}_{n}(t)=b_{n-1}\phi_{n-1}(t)-b_{n}\phi_{n+1}(t)$ with the initial conditions $\phi_{n}(0)=\delta_{n,0}$ The recursion relation suggests that the Lanczos coefficients $b_{n}$ are hopping amplitudes for the initial operator $\mathcal{O}_{0}$ localized at the initial site to explore the Krylov chain. With time, the operator gains support away from the origin in the Krylov chain reflects the growth of complexity as higher Krylov basis vectors are required in operator expansion. To quantify this, one defines the average position of the operator in the Krylov chain — called the Krylov complexity as
@@
where $\mathcal{K}=\sum_{n=0}^{K-1}n|O_{n})(O_{n}|$ is the position operator in the Krylov chain. For our purpose, we will use the infinite-temperature inner product, also known as the Frobenius inner product :
@@ II.2.2 Discrete-time evolution
To formulate the K-complexity for a system with discrete-time evolution such that $\mathcal{O}_{n}=U^{\dagger}_{n}\mathcal{O}_{n-1}U_{n}=\mathcal{U}_{n}[\mathcal{O}_{n}]$, where $n=1,2,\ldots$, $U_{n}$ is the unitary operator which describes the evolution of the system at time-step $t=nT$ with step-size $T$, and $\mathcal{U}_{n}$ is the unitary superoperator given by $\mathcal{U}_{n}=U_{n}^{\dagger}\bullet U_{n}$. We define the Krylov basis by choosing $|O_{0})=\mathcal{O}_{0}$ and then recursively orthogonalizing each $\mathcal{O}_{n}$ with all the $|O_{n})$ for $i<n$. This choice of basis (represented by $\mathbb{O}$) maximizes the cost function defined as [59]
@@
with respect to an arbitrary choice of orthonormal set $\mathbb{B}=\{|B_{i}):i=0,1,2,\ldots\}$. The coefficients $\zeta_{i}$ (referred to as the “weight function”) are positive, increasing the sequence of real numbers. The optimal choice of the basis follows from the induction method. We fix the initial state as the first state of the Krylov basis $\mathcal{O}_{0}=|O_{0})$. Assume the first $N$ vectors of certain basis $\mathbb{B}$ are same as the Krylov basis, i.e., $|B_{i}\rangle=|O_{i}\rangle$ for $i=0,1,\ldots,N-1$. By assumption
@@
Therefore, the cost associated with the two basis are the same for $n\leq N-1$. The next state $|O_{N}\rangle$ into a part belonging to the Krylov subspace $|O_{i}\rangle$, for $i\leq N-1$, and a part perpendicular to it i.e.
@@
Since the evolution operator itself is used as a generator of the Krylov basis, we can bound the maximum possible growth of the K-complexity. For any system evolving unitarily in discrete-time, the K-complexity can grow at most linearly with time $t$. To prove this, we consider maximizing the complexity at anytime $t$ with respect to the expansion coefficients, i.e.,
@@
with constraint $\sum_{i}|\phi_{i,n}|^{2}=1$. From the orthonormalization construction, it follows that the expansion coefficients $\phi_{i,n}=0\ \forall\ i>n$. Therefore, it follows that the complexity can at most grows linearly with time $t$ corresponding to the case where $\phi_{i,n}=\delta_{i,n}$. In terms of the operator, it means that $(O_{n^{\prime}}|O_{n})\ \forall\ n^{\prime}<n$ which corresponds to the maximally ergodic regime universal in chaotic systems [60].
In the case where the evolution operator is time-independent, we can define a quasi-Hamiltonian $H_{q}$ such that $U=e^{-iH_{q}T}$. In this case, the unitary operator $U$ takes upper Hessenberg form in the Krylov basis [60], given by
@@
with $a_{0}=c_{0}$. The quasi-Hamiltonian $H_{q}$ can itself be used in the continuous-time setting to define complexity using the notion discussed in Sec. II.2.1. To investigate the relation between the complexities, we consider the following expansion
@@
where $\mathcal{L}_{q}$ is Liouvillian superoperator associated with the Hamiltonian $H_{q}$. The Taylor series is truncated at order $K_{c}$ which depends on the operator $\mathcal{L}_{q}$, step-size $T$ and error tolerance [61]. From Eq. (12), it follows that the operator at first time-step $\mathcal{O}_{1}$ will be linear combination of operators $\{\mathcal{L}^{k}_{q}\mathcal{O}_{0}\ |\ k=0,1,\ldots,K_{c}\}$. In other words, the first Krylov basis vector
@@
where the set $\mathbb{O}^{(q)}=\{|O^{(q)}_{k})\ |\ k=0,1,\ldots\}$ corresponds to the Krylov basis associated with the Liouvillian $\mathcal{L}_{q}$. In general, we will assume that there exists a cutoff value $K_{c}^{(i)}$ such that
@@ III Discrete-time quantum walks
corresponding to parameters $(0,\theta,0,3\pi/2)$ which previously was studied in the context of the Dirac dynamic [62]. In momentum basis, the unitary operator can be diagonalized to obtain the dispersion relation in space-time continuum limit [63]
@@
It’s important to note that group velocity is maximum (equals to $1$) at $\theta=0$ for all $k$ which corresponds to identity as coin operator. The coin parameter $\theta$ controls the variance $\sigma^{2}$ of the probability distribution in the position space and this distribution spreads quadratically faster ($\sigma^{2}\approx[1-t^{2}\sin\theta])$ in position space when compared to the classical random walk [64].
@@ III.1 Disordered discrete-time quantum walk
The mean group velocity drops to zero faster for a walk with spatial disorder resulting in strong localization compared to temporal disorder which leads to weak localization. the localization length is usually a function of the coin parameter $\theta$ given as $\zeta=[\ln\cos\theta]^{-1}$. Both of these disorders were studied extensively in enhancing the entanglement and non-Markovianity generated between the internal and external degrees of freedom [71, 70].
@@ IV.1 Out-of-time-ordered correlator
In integrable systems such as quantum spin systems [72], one can analytically find the squared commutator to exactly describe the nature of the wavefront and decay. In addition, although, quantum walk is also an integrable model, it turns out to be particularly hard to find out exact form of squared commutator. This is because the initial operators are local in position space, and the inverse fourier transform becomes hard due to the complexity of functions. However, in case of the Hadamard quantum walk ($\theta=\pi/4$), we can use the known asymptotic results of wavefunction to vaguely argue the nature of decay and wavefront [73]. To this end, the asymptotics for the wave function can be described in three regions using $\alpha=l/t$. The wave function is essentially uniformly spread over the interval between $\mp 1/\sqrt{2}=\mp v_{B}$ where its gross behaviour is like $1/\sqrt{t}$. Outside this interval, the wave function dies out much faster than any inverse polynomial in $t$. At the “wavefront” (±$\pm v_{B}$), there are two peaks of width $\mathcal{O}(t^{1/3})$, where the wave function goes as $t^{-1/3}$. If we approximately translate these wave function behaviour to operator $W^{\mu}_{l}(t)$, we conclude that the squared commutator decays as $\sim 1/t^{2}$ in the long-time limit and goes as $t^{-4/3}$ at the wavefront.
@@ IV.2 K-complexity
where the blue dot represent the presence of term $P_{ij}=|i\rangle\langle j|$ and multiple presence of blue dot represents the linear combination of these terms in operator expansion. The inner product between operators $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$ depends on the expansion term corresponding to where dot color matches. The inner product between the operators in Eq. (32) is zero i.e. the operators are orthonormal to each other since there are no common blue dots. At $t=2,3$, the operators are given by